Suppose the risk-free rate is 2.74% and an analyst assumes a market risk premium of 6.11%. Firm A just paid a dividend of $1.03 per share. The analyst estimates the β of Firm A to be 1.44 and estimates the dividend growth rate to be 4.06% forever. Firm A has 277.00 million shares outstanding. Firm B just paid a dividend of $1.63 per share. The analyst estimates the β of Firm B to be 0.88 and believes that dividends will grow at 2.79% forever. Firm B has 182.00 million shares outstanding. What is the value of Firm B?
Answer #1

Since the Firm B just paid the dividend, so

Current Dividend per share of Firm B = D0 = $1.63 , Constant growth rate of dividends of Firm B = g = 2.79%

Now Using CAPM

Required rate of return of Stock of Firm B = r = Risk free rate + Beta x Market risk premium = 2.74% + 0.88 x 6.11% = 2.74% + 5.3768% = 8.1168%

Expected dividend next year = D1 = D0(1+g) = 1.63(1+2.79%) = 1.63 x 1.0279 = 1.6754

Now using constant growth rate model, we get

Price of stock of Firm B = D1 / (r - g) = 1.6754 / (8.1168% - 2.79%) = 1.6754 / 5.3268% = $31.4522

Value of Equity of Firm B = Price of stock of Firm x no of shares outstanding = $31.4522 x 182 million = 5724.3004 million

As debt of Firm is not given, we assume Value of debt of Firm B = 0

Value of Firm B = Value of equity + Value of debt = 5724.3004 + 0 = 5724.3004 million

Hence Value of Firm B = $5724.3004 million